A variable is normally distributed with mean 22 and standard deviation 5. Use your graphing calculator to find each of the following areas. Write your answers in decimal form. Round to the nearest thousandth as needed.

a) Find the area to the left of 24. 
b) Find the area to the left of 18. 
c) Find the area to the right of 19. 
d) Find the area to the right of 25. 
e) Find the area between 18 and 28.
Answer #1

X\sim N(22,~5^2).\\\\ (a)~P(X<23)=0.579\\\\ (b)~P(X<20)=0.345\\\\ (c)~P(X>21)=1-P(X\leq 21)=0.579\\\\ (d)~P(X>27)=1-P(X\leq 27)=0.159\\\\ (e)~P(20\leq X\leq 30)=0.601

answered by: ANURANJAN SARSAM